In a sale there is 25% off all prices.A chair costs £45 in the sale.How much was it before the sale?
Found 2 solutions by TimothyLamb, ewatrrr:
Answer by TimothyLamb(4379)

x = original price
---
sale price:
(1 - 0.25)x = 45
(0.75)x = 45
x = 45/0.75
---
answer:
x = 60
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by ewatrrr(24785)

paying 75%
.75P+=+45
P = 45/.75 = £60